A sample of 100. grams of solid NaHCO3 was placed in a previously evacuated rigid 5.00 Liter container and heated to 160C. Some of the original solid remained and the total pressure in the container was 7.76 atmospheres when equilibrium was reached. Calculate the number of moles of H2O(g) and CO2 (g)present at equilibrium



1
Expert's answer
2021-02-12T05:00:24-0500

First write the balance equation

2Na2Co3= Na2co3+H2O+co2

Ptotal= Pwater+ pco2

We know that pressure is directly proportional to the number of mole according to ideal gas law.

Pwater/pco2= mole of water/mole of co2

Pwater= ptotal/2

Pwater= 7.76/2=3.88atm


Now we know that PV=nRT

3.88×5= n×0.0831×433

n= 0.54mole